OpenStudy (anonymous):

20 is 125% of what number?

OpenStudy (anonymous):

20=1.25x

hero (hero):

is/of = %/100

hero (hero):

20/x = 125/100

hero (hero):

cross multiply: 125x = 2000

OpenStudy (anonymous):

so its 12?

hero (hero):

x = 2000/125 x = 16